OpenStudy (anonymous):

Brackett.. can you help me out with this problem.. Solve the following equation by clearing fractions. show work x/6+2=4/5-x

OpenStudy (anonymous):

is it x 4 ------ + 2 = ---- -x ??????? 6 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then x= -36/35

OpenStudy (anonymous):

do you need the working??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

How do you get rid of the fractions

OpenStudy (anonymous):

x 4 x 4 ------ + 2 = ---- -x = ------ + x = ---- - 2 6 5 6 5 x + 6x 4-10 7x -6 ------ = -------- = ------ = ------ 6 5 6 5 cross multiplying, u get 35x = -36 so x = -36/35 If satisfied, pls click the GOOD ANSWER button :)